OpenStudy (anonymous):

Two investments are made totalling 8800 Pesos. For a certain year, these investments yield 1326 Pesos in simple interest. Part of the 8800 is invested at 16% and part at 14%. Find the amount invested at each rate ???

OpenStudy (anonymous):

x*0.16+(8800-x)*0.14 = 1326 ... solve for x

OpenStudy (gw2011):

Let x=Amount invested at 16% Let y=Amount invested at 14% Now you set up two equations: 0.16x+0.14y=1326 x+y=8800 Multiply this equation by 0.14 to eliminate y and find x 0.16x+0.14y=1326 0.14x+0.14y=1232 Subtract this equation from the first one to find x 0.02x=94 x=4700 invested at 16% x+y=8800 y=8800-x y=8800-4700 y=4100 invested at 14%